A man is 2/5 the way walking accross a railroad bridge whenhe hears a train approaching from behind him. He can reach either end of the bridge just as the train reaches the same point. If the trainis traveling at 60 miles per hour(mph), at what average speed must the man run?
Answer by Paul(988)

2(2/5)=4/5
5/5-4/5=1/5
That means that the man would run 1/5 of the bridge, and it would take the train to run the whole 5/5 of the bridge
1x/5=5/5
x=5
60/5=12
Hence, the speed of the man is 12mph.
Paul.